# Invariants of Finite Orthogonal Groups of Plus Type in Odd Characteristic

### v2

@@ 5.18. Computing the Borel Invariants
Note that $y_{1}^{(q-1)q^{n-2}}$ is the largest monomial of degree $(q-1)q^{n-2}$. The elements of $W$ permute the variables. Every term of $N(y_{1})$ is divisible by $y_{1}$. Therefore the lead term of $N(y_{1})^{q-1}w$ is $y_{1}^{(q-1)q^{n-2}}$ if and only if $w$ stablises $y_{1}$. Since elements of $B$ do not increase the $y_{1}$-degree, the only coset representatives that contribute to the coefficient of $y_{1}^{(q-1)q^{n-2}}$ in $\mathcal{R}(N(y_{1})^{q-1})$ are those of the form $wb$ were $w$ stablises $y_{1}$. If $w$ stabilises $y_{1}$, then all of the coset representatives of the form $wb$ make the same contribution to the coefficient. The number of coset representatives associated to a double coset is either $1$, if $P\cap w^{-1}Pw=P$, or congruent to 00 modulo $p$. The only elements of $W$ for which $P\cap w^{-1}Pw=P$ are $1$ and $\omega$. Therefore the coefficient of $y_{1}^{(q-1)q^{n-2}}$ in $\mathcal{R}(N(y_{1})^{q-1})$ is $(1/2)2=1$. ∎
@@ 6. The Invariants of ${{\text{O}^{+}_{{4}}(\mathbb{F}_{q})}}$
The ring of invariants for ${{\text{O}^{+}_{{4}}(\mathbb{F}_{q})}}$ was computed by Huah Chu [8]. In this section we provide an alternative computation and show that Theorem 4.6 is satisfied for $m=2$. This provides the base case for the inductive proof of the theorem.
The group ${{\text{O}^{+}_{{2}}(\mathbb{F}_{q})}}$ has order $2(q-1)$ and the ring of invariants is generated by $\xi_{0}=y_{1}x_{1}=u_{1}$ and $d_{1,1}=y_{1}^{q-1}+x_{1}^{q-1}=\xi_{1}/\xi_{0}$.
@@
$u_{2}=\xi_{0}^{q}(\xi_{2}+c_{2,2})-\xi_{1}^{q+1}$ with $c_{2,2}\in R_{1}$. Furthermore
$c_{2,2}=\xi_{0}\xi_{1}^{q-1}+\xi_{0}^{q+2}\xi_{1}^{q-3}-\sum_{j=2}^{(q-1)/2}(-1)^{j}\xi_{1}^{q-1-2j}\xi_{0}^{1+j(q+1)}\frac{(q-2-j)!}{j!((q-1)-2j)!}$
and $\ker(\Phi_{2,1})=u_{2}R_{2}$.
@@
Working in $S_{1}$, we have $\xi_{1}=\xi_{0}d_{1,1}$ where $d_{1,1}=y^{q-1}+x^{q-1}$. Applying $\mathcal{P}^{q}$, using the Cartan identity and the stability condition, gives $\xi_{2}=\xi_{1}\mathcal{P}^{q-1}(d_{1,1})+\xi_{0}^{q}\mathcal{P}^{q-2}(d_{1,1})=\xi_{1}d_{1,1}^{q}+\xi_{0}^{q}\mathcal{P}^{q-2}(d_{1,1})$. Multiplying by $\xi_{0}^{q}$ gives $\xi_{0}^{q}\xi_{2}=\xi_{1}^{q+1}+\xi_{0}^{2q}\mathcal{P}^{q-2}(d_{1,1})$. We will show that $\xi_{0}^{q}\mathcal{P}^{q-2}(d_{1,1})\in R_{1}$. From this it, follows that the relation lifts to $S_{2}$, giving a non-zero element of $\ker(\Phi_{2,1})$. Using Lemma 4.2, the element is divisible by $u_{2}$ Comparing lead terms using the lexicographic order in $S_{2}$ gives $u_{2}=\xi_{0}^{q}(\xi_{2}+c_{2,2})-\xi_{1}^{q+1}$ with $c_{2,2}=-\xi_{0}^{q}\mathcal{P}^{q-2}(d_{1,1})$. A simple calculation gives $\mathcal{P}^{q-2}(d_{1,1})=-(y^{q^{2}-2q+1}+x^{q^{2}-2q+1})$. Taking $a=y^{q-1}$ and $b=x^{q-1}$, we have $-\mathcal{P}^{q-2}(d_{1,1})=a^{q-1}+b^{q-1}$. We can write the symmetric polynomial $a^{q-1}+b^{q-1}$ as a polynomial in $\sigma_{1}=a+b$ and $\sigma_{2}=ab$. A closed form for this expression over the integers is
@@
see [16] formula (1) or (10) (note there is a typographical error in [16](10) where $i+j+1$ is written but $i+j-1$ is intended). Observe that $\sigma_{2}=\xi_{0}^{q-1}$, $\sigma_{1}=d_{1,1}$ and $\xi_{0}\sigma_{1}=\xi_{1}$. Multiplying the above expression by $\xi_{0}^{q}$ and substituting gives
$\xi_{0}^{q}\mathcal{P}^{q-2}(d_{1,1})=\sum_{j=0}^{(q-1)/2}(-1)^{j}\xi_{1}^{q-1-2j}\xi_{0}^{1+j(q+1)}\frac{(q-2-j)!}{j!((q-1)-2j)!}\in R_{1},$
@@
$=\frac{(2j)!}{j!(j+1)!}=\frac{1}{j+1}\binom{2j}{j}\ .\qed$
Note that $u_{2}=(\xi_{2}+c_{2,2})u_{1}^{q}-\xi_{1}(u_{1}d_{1,1})^{q}$. Therefore the displayed equation in part(a) of Theorem 4.6 holds for $m=2$.
@@
(a) $\mathcal{P}^{1}(c_{2,2})=\xi_{1}^{q}$.
(b) $\mathcal{P}^{i}(c_{2,2})=0$ for $1<i<q$.
(c) $\mathcal{P}^{q^{2}}(c_{2,2})\equiv_{\langle\xi_{0}^{q^{2}}\rangle}\xi_{1}^{q^{2}-q+1}-\xi_{0}^{q}\xi_{1}^{q^{2}-2q}\xi_{2}$.
(d) $\mathcal{P}^{q^{2}-q}(c_{2,2})\equiv_{\langle\xi_{0}^{q^{2}}\rangle}\xi_{0}\xi_{2}^{q-1}-\xi_{0}^{q}\xi_{1}^{q^{2}-2q+1}+\xi_{0}^{2q}\xi_{1}^{q^{2}-3q}\xi_{2}$.
@@
Since $\mathcal{P}^{1}(u_{2})=0$, we have
$0=\mathcal{P}^{1}(\xi_{0}^{q}(\xi_{2}+c_{2,2})-\xi_{1}^{q+1})=\xi_{0}^{q}(\xi_{1}^{q}+\mathcal{P}^{1}(c_{2,2}))-2\xi_{0}^{q}\xi_{1}^{q}.$
Therefore $\mathcal{P}^{1}(c_{2,2})=\xi_{1}^{q}$, proving (a). For $1<i<q$, we have
$0=\mathcal{P}^{i}(\xi_{0}^{q}(\xi_{2}+c_{2,2})-\xi_{1}^{q+1})=\xi_{0}^{q}\mathcal{P}^{i}(c_{2,2}),$
giving $\mathcal{P}^{i}(c_{2,2})=0$. Since $c_{2,2}\equiv_{\langle\xi_{0}^{q}\rangle}\xi_{0}\xi_{1}^{q-1}$, it follows from Lemma 3.6 that
$\mathcal{P}^{q^{2}}(c_{2,2})\equiv_{\langle\xi_{0}^{q^{2}}\rangle}\mathcal{P}^{q^{2}}(\xi_{0}\xi_{1}^{q-1})=\xi_{1}^{q^{2}-q+1}-\xi_{0}^{q}\xi_{1}^{q^{2}-2q}\xi_{2},$
@@
$\mathcal{P}^{q^{2}-q}(\xi_{0}\xi_{1}^{q-1})=\xi_{0}\mathcal{P}^{q^{2}-q}(\xi_{1}^{q-1})+\xi_{1}\mathcal{P}^{q^{2}-q-1}(\xi_{1}^{q-1})+\xi_{0}^{q}\mathcal{P}^{q^{2}-q-2}(\xi_{1}^{q-1}).$
Note that $\mathcal{P}^{q^{2}-q}(\xi_{1}^{q-1})=\xi_{2}^{q-1}$, $\mathcal{P}^{q^{2}-q-1}(\xi_{1}^{q-1})=-2\xi_{0}^{q}\xi_{1}^{q^{2}-2q}$ and
$\mathcal{P}^{q^{2}-q-2}(\xi_{1}^{q-1})=-\xi_{1}^{q^{2}-2q+1}-\xi_{2}\mathcal{P}^{q^{2}-2q-2}(\xi_{1}^{q-2})=-\xi_{1}^{q^{2}-q+q}+4\xi_{2}\xi_{0}^{q}\xi_{1}^{q^{2}-3q}.$
Therefore $\mathcal{P}^{q^{2}-q}(\xi_{0}\xi_{1}^{q-1})=\xi_{0}\xi_{2}^{q-1}-3\xi_{0}^{q}\xi_{1}^{q^{2}-2q+1}+4\xi_{0}^{2q}\xi_{1}^{q^{2}-3q}\xi_{2}$. Observe that $\mathcal{P}^{q^{2}-q}(\xi_{0}^{q+2}\xi_{1}^{q-3})\equiv_{\langle\xi_{0}^{q^{2}}\rangle}\xi_{1}^{q}\mathcal{P}^{q^{2}-2q}(\xi_{0}^{2}\xi_{1}^{q-3})$ and
$\mathcal{P}^{q^{2}-2q}(\xi_{0}^{2}\xi_{1}^{q-3})=2\xi_{0}^{q}\xi_{1}^{q^{2}-3q+1}-3\xi_{0}^{2q}\xi_{1}^{q^{2}-4q}\xi_{2}.$
Therefore $\mathcal{P}^{q^{2}-q}(\xi_{0}^{q+2}\xi_{1}^{q-3})\equiv_{\langle\xi_{0}^{q^{2}}\rangle}2\xi_{0}^{q}\xi_{1}^{q^{2}-2q+1}-3\xi_{0}^{2q}\xi_{1}^{q^{2}-3q}$. Suppose $b$ is homogeneous of degree $q^{2}-4q+1$. Then $\mathcal{P}^{q^{2}-q}(\xi_{0}^{2q}b)\equiv_{\langle\xi_{0}^{q^{2}}\rangle}\xi_{1}^{q}\mathcal{P}^{q^{2}-2q}(\xi_{0}^{q}b)$ and, using Lemma 3.6, $\mathcal{P}^{q^{2}-2q}(\xi_{0}^{q}b)=\xi_{0}^{q^{2}}\mathcal{P}^{q^{2}-4q}(b)$. Therefore $\mathcal{P}^{q^{2}-q}(\xi_{0}^{2q}b)\equiv_{\langle\xi_{0}^{q^{2}}\rangle}0$ and
$\mathcal{P}^{q^{2}-q}(c_{2,2})$ $\equiv_{\langle\xi_{0}^{q^{2}}\rangle}$ $\mathcal{P}^{q^{2}-q}(\xi_{0}\xi_{1}^{q-1})+\mathcal{P}^{q^{2}-q}(\xi_{0}^{q+2}\xi_{1}^{q-3})$
$\equiv_{\langle\xi_{0}^{q^{2}}\rangle}$ $\xi_{0}\xi_{2}^{q-1}-\xi_{0}^{q}\xi_{1}^{q^{2}-2q+1}+\xi_{0}^{2q}\xi_{1}^{q^{2}-3q}\xi_{2}$
@@
(a) $u_{2}d_{1,2}-\xi_{0}^{q}\xi_{3}\in R_{2}$ and
$u_{2}d_{1,2}\equiv_{\langle\xi_{0}^{q^{2}}\rangle}\xi_{0}^{q}\xi_{3}-\xi_{1}\xi_{2}^{q}+\xi_{0}\xi_{1}^{q}\xi_{2}^{q-1}.$
(b) $u_{2}d_{2,2}-\xi_{1}^{q}\xi_{3}\in R_{2}$ and
$u_{2}d_{2,2}\equiv_{\langle\xi_{0}^{q^{2}}\rangle}\xi_{1}^{q}\xi_{3}-\xi_{2}^{q+1}-\xi_{0}^{q}\xi_{1}^{q^{2}-q}\xi_{2}.$
@@
Using Lemma 6.1, Lemma 3.5 and the Cartan identity,
$\mathcal{P}^{q^{2}}(u_{2})=\xi_{0}^{q}(\xi_{3}+\mathcal{P}^{q^{2}}(c_{2,2}))-\xi_{1}\xi_{2}^{q}+\xi_{1}^{q}\mathcal{P}^{q^{2}-q}(c_{2,2})+\xi_{0}^{q^{2}}\mathcal{P}^{q^{2}-2q}(c_{2,2}).$
Since $c_{2,2}\in R_{1}$, we have $\mathcal{P}^{i}(c_{2,2})\in R_{2}$ and $u_{2}d_{1,2}-\xi_{0}^{q}\xi_{3}\in R_{2}$. Using Lemma 6.3, we have
$u_{2}d_{1,2}=\mathcal{P}^{q^{2}}(u_{2})$ $\equiv_{\langle\xi_{0}^{q^{2}}\rangle}$ $\xi_{0}^{q}(\xi_{3}+\xi_{1}^{q^{2}-q+1}-\xi_{0}^{q}\xi_{1}^{q^{2}-2q}\xi_{2})-\xi_{1}\xi_{2}^{q}$
$+\xi_{1}^{q}(\xi_{0}\xi_{2}^{q-1}-\xi_{0}^{q}\xi_{1}^{q^{2}-2q+1}+\xi_{0}^{2q}\xi_{1}^{q^{2}-3q}\xi_{2})$
$\equiv_{\langle\xi_{0}^{q^{2}}\rangle}$ $\xi_{0}^{q}\xi_{3}-\xi_{1}\xi_{2}^{q}+\xi_{0}\xi_{1}^{q}\xi_{2}^{q-1}$
@@
$\mathcal{P}^{q^{2}+q}(u_{2})=\xi_{1}^{q}(\xi_{3}+\mathcal{P}^{q^{2}}(c_{2,2}))-\xi_{2}^{q+1}-\xi_{1}^{q^{2}+1}+\xi_{0}^{q^{2}}\mathcal{P}^{q^{2}-q}(c_{2,2}).$
@@
$\mathcal{P}^{q^{2}+q}(u_{2})\equiv_{\langle\xi_{0}^{q^{2}}\rangle}\xi_{1}^{q}\xi_{3}-\xi_{2}^{q+1}-\xi_{1}^{q^{2}+1}+\xi_{1}^{q}(\xi_{1}^{q^{2}-q+1}-\xi_{0}^{q}\xi_{1}^{q^{2}-2q}\xi_{2}).$
Therefore $u_{2}d_{2,2}\equiv_{\langle\xi_{0}^{q^{2}}\rangle}\xi_{1}^{q}\xi_{3}-\xi_{2}^{q+1}-\xi_{0}^{q}\xi_{1}^{q^{2}-q}\xi_{2}$, as required. ∎
@@
(a) $\mathcal{P}^{i}(u_{2})=0$ for $0<i<q$.
(b) $\mathcal{P}^{i}(u_{2})/u_{2}\in R_{2}$ for $i\not\in\{q^{2},q^{2}+q,q^{2}+2q\}$.

### v3

@@ 5.20. Computing the Borel Invariants
Note that $y_{m}^{(q-1)q^{n-2}}$ is the largest monomial of degree $(q-1)q^{n-2}$. The elements of $W$ permute the variables. Every term of $N(y_{m})$ is divisible by $y_{m}$. Therefore the lead term of $N(y_{m})^{q-1}w$ is $y_{m}^{(q-1)q^{n-2}}$ if and only if $w$ stablises $y_{m}$. Since elements of $B$ do not increase the $y_{m}$-degree, the only coset representatives that contribute to the coefficient of $y_{m}^{(q-1)q^{n-2}}$ in $\mathcal{R}(N(y_{m})^{q-1})$ are those of the form $wb$ were $w$ stablises $y_{m}$. If $w$ stabilises $y_{m}$, then all of the coset representatives of the form $wb$ make the same contribution to the coefficient. The number of coset representatives associated to a double coset is either $1$, if $P\cap w^{-1}Pw=P$, or congruent to 0 modulo $p$. The only elements of $W$ for which $P\cap w^{-1}Pw=P$ are $1$ and $\omega$. Therefore the coefficient of $y_{m}^{(q-1)q^{n-2}}$ in $\mathcal{R}(N(y_{m})^{q-1})$ is $(1/2)2=1$. ∎
@@ 6. The Invariants of $\text{O}^{+}_{{4}}(\mathbb{F}_{q})$}}
The ring of invariants for $\text{O}^{+}_{{4}}(\mathbb{F}_{q})$}} was computed by Huah Chu [8]. In this section we provide an alternative computation and show that Lemma 4.6 is satisfied for $m=2$. This provides the base case for the inductive proof of the lemma and implies that Theorem 4.8 holds for $m=2$.
The group $\text{O}^{+}_{{2}}(\mathbb{F}_{q})$}} has order 2(q1)2(q-1) and the ring of invariants is generated by $\xi_{1}=y_{1}x_{1}=u_{1}$ and $d_{1,1}=y_{1}^{q-1}+x_{1}^{q-1}=\xi_{2}/\xi_{1}$.
@@
$u_{2}=\xi_{1}^{q}(\xi_{3}+c_{3,2})-\xi_{2}^{q+1}$ with $c_{3,2}\in R_{2}$. Furthermore
$c_{3,2}=\xi_{1}\xi_{2}^{q-1}+\xi_{1}^{q+2}\xi_{2}^{q-3}-\sum_{j=2}^{(q-1)/2}(-1)^{j}\xi_{2}^{q-1-2j}\xi_{1}^{1+j(q+1)}\frac{(q-2-j)!}{j!((q-1)-2j)!}$
and $\ker(\Phi_{3,1})=u_{2}R_{3}$.
@@
Working in $S_{1}$, we have $\xi_{2}=\xi_{1}d_{1,1}$ where $d_{1,1}=y^{q-1}+x^{q-1}$. Applying $\mathcal{P}^{q}$, using the Cartan identity and the stability condition, gives $\xi_{3}=\xi_{2}\mathcal{P}^{q-1}(d_{1,1})+\xi_{1}^{q}\mathcal{P}^{q-2}(d_{1,1})=\xi_{2}d_{1,1}^{q}+\xi_{1}^{q}\mathcal{P}^{q-2}(d_{1,1})$. Multiplying by $\xi_{1}^{q}$ gives $\xi_{1}^{q}\xi_{3}=\xi_{2}^{q+1}+\xi_{1}^{2q}\mathcal{P}^{q-2}(d_{1,1})$. We will show that $\xi_{1}^{q}\mathcal{P}^{q-2}(d_{1,1})\in R_{2}$. From this it, follows that the relation lifts to $S_{2}$, giving a non-zero element of $\ker(\Phi_{3,1})$. Using Lemma 4.2, the element is divisible by $u_{2}$ Comparing lead terms using the lexicographic order in $S_{2}$ gives $u_{2}=\xi_{1}^{q}(\xi_{3}+c_{3,2})-\xi_{2}^{q+1}$ with $c_{3,2}=-\xi_{1}^{q}\mathcal{P}^{q-2}(d_{1,1})$. A simple calculation gives $\mathcal{P}^{q-2}(d_{1,1})=-(y^{q^{2}-2q+1}+x^{q^{2}-2q+1})$. Taking $a=y^{q-1}$ and $b=x^{q-1}$, we have $\mathcal{P}^{q-2}(d_{1,1})=a^{q-1}+b^{q-1}$. We can write the symmetric polynomial $a^{q-1}+b^{q-1}$ as a polynomial in $\sigma_{1}=a+b$ and $\sigma_{2}=ab$. A closed form for this expression over the integers is
@@
see [17] formula (1) or (10) (note there is a typographical error in [17](10) where $i+j+1$ is written but i+j1i+j-1 is intended). Observe that $\sigma_{2}=\xi_{1}^{q-1}$, $\sigma_{1}=d_{1,1}$ and $\xi_{1}\sigma_{1}=\xi_{2}$. Multiplying the above expression by $\xi_{1}^{q}$ and substituting gives
$\xi_{1}^{q}\mathcal{P}^{q-2}(d_{1,1})=\sum_{j=0}^{(q-1)/2}(-1)^{j}\xi_{2}^{q-1-2j}\xi_{1}^{1+j(q+1)}\frac{(q-2-j)!}{j!((q-1)-2j)!}\in R_{2},$
@@
$=\frac{(2j)!}{j!(j+1)!}=\frac{1}{j+1}\binom{2j}{j}\$$\qed$
@@
If $u_{2}$ divides $f$ in $S_{2}$ and $f\in R_{3}$ then $f/u_{2}\in R_{3}$.
@@
From Lemma 6.1, we have $\ker(\Phi_{3,1})=u_{2}R_{3}$. The result follows using Lemma 4.2(b) and the argument given in Remark 4.7. ∎
Lemma 6.4.
(a) $\mathcal{P}^{i}(u_{2})=0$ for $0<i<q$.
(b) $\mathcal{P}^{i}(u_{2})/u_{2}\in R_{3}$ for $i\not\in\{q^{2},q^{2}+q,q^{2}+2q\}$.
Proof.
As an element of $S_{2}$, $u_{2}$ is a product of pairwise relatively prime linear forms. Therefore, using Lemma 3.5, $u_{2}$ divides $\mathcal{P}^{i}(u_{2})$ in $S_{2}$ for all $i$. If $\mathcal{P}^{i}(u_{2})\in R_{3}$, it follows from Lemma 6.3 that $\mathcal{P}^{i}(u_{2})/u_{2}\in R_{3}$. Recall from Lemma 6.1 that $u_{2}=\xi_{1}^{q}(\xi_{3}+c_{3,2})-\xi_{2}^{q+1}$ with $c_{3,2}\in R_{2}$. Therefore, using the Cartan identity, Lemma 3.4, and Corollary 3.6, we have $\mathcal{P}^{i}(u_{2})\in R_{2}\subset R_{3}$ for $0<i<q$. Since $u_{2}$ is linear in $\xi_{3}$, this give $\mathcal{P}^{i}(u_{2})=0$ for $0<i<q$, proving (a). Again using Corollary 3.6, for $i\not\in\{q^{2},q^{2}+q,q^{2}+2q\}$, we have $\mathcal{P}^{i}(u_{2})\in R_{3}$. Therefore, if $i\not\in\{q^{2},q^{2}+q,q^{2}+2q\}$, we have $\mathcal{P}^{i}(u_{2})/u_{2}\in R_{3}$, proving (b). ∎
Lemma 6.5.
(a) $\mathcal{P}^{1}(c_{3,2})=\xi_{2}^{q}$.
(b) $\mathcal{P}^{i}(c_{3,2})=0$ for $1<i<q$.
(c) $\mathcal{P}^{q^{2}}(c_{3,2})\equiv_{\langle\xi_{1}^{q^{2}}\rangle}\xi_{2}^{q^{2}-q+1}-\xi_{1}^{q}\xi_{2}^{q^{2}-2q}\xi_{3}$.
(d) $\mathcal{P}^{q^{2}-q}(c_{3,2})\equiv_{\langle\xi_{1}^{q^{2}}\rangle}\xi_{1}\xi_{3}^{q-1}-\xi_{1}^{q}\xi_{2}^{q^{2}-2q+1}+\xi_{1}^{2q}\xi_{2}^{q^{2}-3q}\xi_{3}$.
Proof.
Using Lemma 6.4, we have $\mathcal{P}^{1}(u_{2})=0$. Thus
$0=\mathcal{P}^{1}(\xi_{1}^{q}(\xi_{3}+c_{3,2})-\xi_{2}^{q+1})=\xi_{1}^{q}(\xi_{2}^{q}+\mathcal{P}^{1}(c_{3,2}))-2\xi_{1}^{q}\xi_{2}^{q}.$
Therefore $\mathcal{P}^{1}(c_{3,2})=\xi_{2}^{q}$, proving (a). Again using Lemma 6.4, for $1<i<q$, we have
$0=\mathcal{P}^{i}(u_{2})=\mathcal{P}^{i}(\xi_{1}^{q}(\xi_{3}+c_{3,2})-\xi_{2}^{q+1})=\xi_{1}^{q}\mathcal{P}^{i}(c_{3,2}),$
giving $\mathcal{P}^{i}(c_{3,2})=0$. Since $c_{3,2}\equiv_{\langle\xi_{1}^{q}\rangle}\xi_{1}\xi_{2}^{q-1}$, it follows from Lemma 3.7 that
$\mathcal{P}^{q^{2}}(c_{3,2})\equiv_{\langle\xi_{1}^{q^{2}}\rangle}\mathcal{P}^{q^{2}}(\xi_{1}\xi_{2}^{q-1})=\xi_{2}^{q^{2}-q+1}-\xi_{1}^{q}\xi_{2}^{q^{2}-2q}\xi_{3},$
@@
$\mathcal{P}^{q^{2}-q}(\xi_{1}\xi_{2}^{q-1})=\xi_{1}\mathcal{P}^{q^{2}-q}(\xi_{2}^{q-1})+\xi_{2}\mathcal{P}^{q^{2}-q-1}(\xi_{2}^{q-1})+\xi_{1}^{q}\mathcal{P}^{q^{2}-q-2}(\xi_{2}^{q-1}).$
Note that $\mathcal{P}^{q^{2}-q}(\xi_{2}^{q-1})=\xi_{3}^{q-1}$, $\mathcal{P}^{q^{2}-q-1}(\xi_{2}^{q-1})=-2\xi_{1}^{q}\xi_{2}^{q^{2}-2q}$ and
$\mathcal{P}^{q^{2}-q-2}(\xi_{2}^{q-1})=-\xi_{2}^{q^{2}-2q+1}-\xi_{3}\mathcal{P}^{q^{2}-2q-2}(\xi_{2}^{q-2})=-\xi_{2}^{q^{2}-2q+1}+4\xi_{3}\xi_{1}^{q}\xi_{2}^{q^{2}-3q}.$
Therefore $\mathcal{P}^{q^{2}-q}(\xi_{1}\xi_{2}^{q-1})=\xi_{1}\xi_{3}^{q-1}-3\xi_{1}^{q}\xi_{2}^{q^{2}-2q+1}+4\xi_{1}^{2q}\xi_{2}^{q^{2}-3q}\xi_{3}$. Observe that $\mathcal{P}^{q^{2}-q}(\xi_{1}^{q+2}\xi_{2}^{q-3})\equiv_{\langle\xi_{1}^{q^{2}}\rangle}\xi_{2}^{q}\mathcal{P}^{q^{2}-2q}(\xi_{1}^{2}\xi_{2}^{q-3})$ and
$\mathcal{P}^{q^{2}-2q}(\xi_{1}^{2}\xi_{2}^{q-3})=2\xi_{1}^{q}\xi_{2}^{q^{2}-3q+1}-3\xi_{1}^{2q}\xi_{2}^{q^{2}-4q}\xi_{3}.$
Therefore $\mathcal{P}^{q^{2}-q}(\xi_{1}^{q+2}\xi_{2}^{q-3})\equiv_{\langle\xi_{1}^{q^{2}}\rangle}2\xi_{1}^{q}\xi_{2}^{q^{2}-2q+1}-3\xi_{1}^{2q}\xi_{2}^{q^{2}-3q}\xi_{3}$. Suppose $b$ is homogeneous of degree $q^{2}-4q+1$. Then $\mathcal{P}^{q^{2}-q}(\xi_{1}^{2q}b)\equiv_{\langle\xi_{1}^{q^{2}}\rangle}\xi_{2}^{q}\mathcal{P}^{q^{2}-2q}(\xi_{1}^{q}b)$ and, using Lemma 3.7, $\mathcal{P}^{q^{2}-2q}(\xi_{1}^{q}b)=\xi_{1}^{q^{2}}\mathcal{P}^{q^{2}-4q}(b)$. Therefore $\mathcal{P}^{q^{2}-q}(\xi_{1}^{2q}b)\equiv_{\langle\xi_{1}^{q^{2}}\rangle}0$ and
$\mathcal{P}^{q^{2}-q}(c_{3,2})$ $\equiv_{\langle\xi_{1}^{q^{2}}\rangle}$ $\mathcal{P}^{q^{2}-q}(\xi_{1}\xi_{2}^{q-1})+\mathcal{P}^{q^{2}-q}(\xi_{1}^{q+2}\xi_{2}^{q-3})$
$\equiv_{\langle\xi_{1}^{q^{2}}\rangle}$ $\xi_{1}\xi_{3}^{q-1}-\xi_{1}^{q}\xi_{2}^{q^{2}-2q+1}+\xi_{1}^{2q}\xi_{2}^{q^{2}-3q}\xi_{3}$
@@
(a) $u_{2}d_{1,2}-\xi_{1}^{q}\xi_{4}\in R_{3}$ and
$u_{2}d_{1,2}\equiv_{\langle\xi_{1}^{q^{2}}\rangle}\xi_{1}^{q}\xi_{4}-\xi_{2}\xi_{3}^{q}+\xi_{1}\xi_{2}^{q}\xi_{3}^{q-1}.$
(b) $u_{2}d_{2,2}-\xi_{2}^{q}\xi_{4}\in R_{3}$ and
$u_{2}d_{2,2}\equiv_{\langle\xi_{1}^{q^{2}}\rangle}\xi_{2}^{q}\xi_{4}-\xi_{3}^{q+1}-\xi_{1}^{q}\xi_{2}^{q^{2}-q}\xi_{3}.$
@@
Using Lemma 6.1, Corollary 3.6 and the Cartan identity,
$\mathcal{P}^{q^{2}}(u_{2})=\xi_{1}^{q}(\xi_{4}+\mathcal{P}^{q^{2}}(c_{3,2}))-\xi_{2}\xi_{3}^{q}+\xi_{2}^{q}\mathcal{P}^{q^{2}-q}(c_{3,2})+\xi_{1}^{q^{2}}\mathcal{P}^{q^{2}-2q}(c_{3,2}).$
Since $c_{3,2}\in R_{2}$, we have $\mathcal{P}^{i}(c_{3,2})\in R_{3}$ and $u_{2}d_{1,2}-\xi_{1}^{q}\xi_{4}\in R_{3}$. Using Lemma 6.5, we have
$u_{2}d_{1,2}=\mathcal{P}^{q^{2}}(u_{2})$ $\equiv_{\langle\xi_{1}^{q^{2}}\rangle}$ $\xi_{1}^{q}(\xi_{4}+\xi_{2}^{q^{2}-q+1}-\xi_{1}^{q}\xi_{2}^{q^{2}-2q}\xi_{3})-\xi_{2}\xi_{3}^{q}$
$+\xi_{2}^{q}(\xi_{1}\xi_{3}^{q-1}-\xi_{1}^{q}\xi_{2}^{q^{2}-2q+1}+\xi_{1}^{2q}\xi_{2}^{q^{2}-3q}\xi_{3})$
$\equiv_{\langle\xi_{1}^{q^{2}}\rangle}$ $\xi_{1}^{q}\xi_{4}-\xi_{2}\xi_{3}^{q}+\xi_{1}\xi_{2}^{q}\xi_{3}^{q-1}$
@@
$\mathcal{P}^{q^{2}+q}(u_{2})=\xi_{2}^{q}(\xi_{4}+\mathcal{P}^{q^{2}}(c_{3,2}))-\xi_{3}^{q+1}-\xi_{2}^{q^{2}+1}+\xi_{1}^{q^{2}}\mathcal{P}^{q^{2}-q}(c_{3,2}).$
@@
$\mathcal{P}^{q^{2}+q}(u_{2})\equiv_{\langle\xi_{1}^{q^{2}}\rangle}\xi_{2}^{q}\xi_{4}-\xi_{3}^{q+1}-\xi_{2}^{q^{2}+1}+\xi_{2}^{q}(\xi_{2}^{q^{2}-q+1}-\xi_{1}^{q}\xi_{2}^{q^{2}-2q}\xi_{3}).$
Therefore $u_{2}d_{2,2}\equiv_{\langle\xi_{1}^{q^{2}}\rangle}\xi_{2}^{q}\xi_{4}-\xi_{3}^{q+1}-\xi_{1}^{q}\xi_{2}^{q^{2}-q}\xi_{3}$, as required. ∎
@@
(a) For $0<i<q$, $\mathcal{P}^{i}(d_{1,2})$ and $\mathcal{P}^{i}(d_{2,2})$ lie in $R_{3}$.
(b) For $0<i<q^{3}$, write $i=\ell q^{2}+r$ with $r<q^{2}$. Then $u_{2}^{\ell+1}\mathcal{P}^{i}(d_{1,2})$ and $u_{2}^{\ell+1}\mathcal{P}^{i}(d_{2,2})$ lie in $R_{4}$ with $\xi_{4}$-degree at most $\ell+1$.